OpenStudy (anonymous):

32 + (8 - 2) • 4 - 6/3

OpenStudy (anonymous):

its not 32, its 3 squared

OpenStudy (anonymous):

your answer is 31 . we'll first solve the bracket

OpenStudy (anonymous):

hint.. \[\Large 32+(8-2)\cdot 4 -6\sqrt3=32+6\cdot 4-6\sqrt3\]

OpenStudy (anonymous):

apply BODMAS i.e Bracket Of Division Mulitiplication Addition Substraction... z answer is 25

OpenStudy (anonymous):

@nataliaviruet you said number three is under square root didn't you ?

OpenStudy (anonymous):

ahh ok its 9 not 32 zen \[9+\subset6\supset \times4-2\] zen simplify to get 31

OpenStudy (anonymous):

I guess I got it wrong. O_O didn't I ?

OpenStudy (anonymous):

lol its 31

OpenStudy (anonymous):

@nataliaviruet you said: nataliaviruet This is the asker its not 32, its 3 squared I see it like: 32 + (8 - 2) • 4 - 6/3 and I thought it must be : \[\Large 32+(8-2)\cdot 4-\frac63\] but then according to you I put number three under square root and wrote it like: \[\Large 32+(8-2)\cdot 4-6\sqrt3\] but now I guess 32 isn't right or what ??